OpenStudy (anonymous):

HELPPPP :( The goals against average (A) for a professional hockey goalie is determined using the formula A = 60 ( g/t) In the formula, g represents the number of goals scored against the goalie and t represents the time played, in minutes. Which is an equivalent equation solved for g?

OpenStudy (danjs):

A =\[A = \frac{ 60*g }{ t }\]

OpenStudy (danjs):

multiply both sides by t, then divide by 60 on both sides

OpenStudy (danjs):

t*A = 60*g just divide by 60 g = ...

OpenStudy (anonymous):

Thank you!! :)